# Photovoltaic power forecasting using quantum machine learning

### v2

@@ Abstract
Predicting solar panel power output is crucial for advancing the transition to renewable energy but is complicated by the variable and non-linear nature of solar energy. This is influenced by numerous meteorological factors, geographical positioning, and photovoltaic cell properties, posing significant challenges to forecasting accuracy and grid stability. Our study introduces a suite of solutions centered around hybrid quantum neural networks designed to tackle these complexities. The first proposed model, the Hybrid Quantum Long Short-Term Memory, surpasses all tested models by achieving mean absolute errors and mean squared errors that are more than $40\%$ lower. The second proposed model, the Hybrid Quantum Sequence-to-Sequence neural network, once trained, predicts photovoltaic power with $16\%$ lower mean absolute error for arbitrary time intervals without the need for prior meteorological data, highlighting its versatility. Moreover, our hybrid models perform better even when trained on limited datasets, underlining their potential utility in data-scarce scenarios. These findings represent progress towards resolving time series prediction challenges in energy forecasting through hybrid quantum models, showcasing the transformative potential of quantum machine learning in catalyzing the renewable energy transition.
quantum machine learning, hybrid quantum neural network, quantum depth infused layer, quantum LSTM, solar energy, photovoltaic power, time series
@@ I Introduction
Electricity generation prediction, especially for photovoltaic (PV) systems, is a crucial tool for renewable energy adoption sabri2022novel ; li2020review . The global economy must radically reduce emissions to stay within the $1.5\degree$C pathway (Paris Agreement) and the transition to renewable energy sources is necessary to achieve these objectives ellabban2014renewable . According to the International Energy Agency, solar PV’s installed power capacity is poised to surpass that of coal by 2027, becoming the largest in the world.
@@
Accurate PV power forecasts are vital for multiple facets of the energy industry such as long-term investment planning, regulatory compliance for avoiding penalties, and renewable energy management across storage, transmission, and distribution activities. Several studies show that an increase in forecasting accuracy reduces electricity generation from conventional sources. Increased accuracy also reduces operating costs of systems through reducing the uncertainty of PV power generation anido2016value . They support improving the stability and sustainability of the power grid through optimizing power flow and counteracting solar power’s intermittent nature lee2021pv . Such predictions are foundational in increasing the economic viability and improving the adoption of solar energy as they inform pricing and economic dispatch strategies, bolster competitiveness and over time reduce reliance on reserve power. Additionally, they assist in managing energy storage effectively and integrating PV systems into the power grid bozorgavari2019two , which is essential for the enduring success of renewable energy solutions diagne2013review .
Traditional methods for predicting PV power have primarily relied on statistical models, machine learning algorithms, or a blend of both ahmed2019review . These approaches encompass a diverse toolkit, ranging from time series forecasting and artificial neural networks sabri2022novel ; gomez2020photovoltaic ; kaloop2021novel , to support vector machines kulkarni2018output ; shi2012forecasting , k-nearest neighbor methods rocha2022computational , and random forest models inman2013solar . However, the intermittent and non-linear nature of solar power generation, influenced by a wide range of meteorological factors, poses a significant challenge to the performance of these conventional models blaga2019current .
@@
In light of these challenges, quantum machine learning (QML) emerges as a promising avenue. This rapidly evolving field, which melds the principles of quantum mechanics with classical machine learning Schuld_2014 ; biamonte2017quantum ; Dunjko.Briegel.2018 ; qml_review_2023 , can offer enhanced capabilities for improving the forecasting accuracy of time series tasks emmanoulopoulos2022quantum , including PV power generation sushmit2023forecasting . QML’s potential arises from quantum features like superposition and entanglement, promising exponential speedups in certain tasks schuld2018implementing . Moreover, QML algorithms produce inherently probabilistic results, aptly suited for prediction tasks, and they may potentially function within an exponentially larger search space, amplifying their efficacy Lloyd2013 ; 365700 ; Lloyd1996UniversalQS ; schuld2020circuit . Nonetheless, implementing quantum algorithms bears its own set of challenges, such as the need for error correction and sensitivity to external interference Aaronson2015 . Yet, in spite of these challenges, hybrid quantum-classical models, especially hybrid quantum neural networks (HQNNs), have showcased their potential in diverse industrial realms, including healthcare jain2022hybrid ; pharma ; steatosis_sagingalieva , energy kurkin2023forecasting ; sushmit2023forecasting , aerospace rainjonneau2023quantum , logistics haboury2023supervised and automotive sagingalieva2023hyperparameter industries.
@@
Figure 1: The input to a hybrid quantum model is presented as a chronological data table, documenting hourly meteorological parameters including ambient temperature ($T_{a}$), module temperature ($T_{m}$), and solar irradiance ($I_{3}$, $I_{15}$), alongside the mean PV power output ($P$). The model is designed to leverage this data to generate predictions of PV power output for a short-term forecast aimed at near-future output, typically the next hour, and a long-term forecast that extends to a broader temporal horizon.
@@
In this article, we present three types of hybrid quantum models as potential solutions for PV power forecasting. We assess the performance of our proposed models using a publicly accessible dataset, encompassing a comprehensive array of meteorological variables as well as hourly mean PV power measurements spanning a $21$-month period. This dataset, along with the data preprocessing and analytical methods employed, is described in detail in Section II.1.
@@
Remarkably, despite having fewer parameters, our hybrid quantum models outperform their classical counterparts in terms of more accurate predictions, including when trained on a reduced dataset. We summarize our conclusions and outline future research directions in Section III.
II Results
@@ II.1 Dataset
To investigate the advantages of hybrid quantum models, we selected a publicly accessible dataset malvoni2016data from a conventional generation plant situated in the Mediterranean region. This dataset not only provides comprehensive data but also enables benchmarking with results from various algorithms available in literature. A comparative analysis of our model’s predictions and those from the study by kaloop2021novel is provided in Section II.5. The dataset, presented as a numerical table showcased in Fig. 1, includes variables like: hourly mean ambient temperature, $T_{a}$; hourly mean module temperature, $T_{m}$; hourly mean solar irradiance recorded on two tilted planes with tilt angles of 3 and 15 degrees, $I_{3}$ and $I_{15}$; and hourly mean PV power, $P$, spanning $21$ months, accounting for more than $500$ days.
Beyond the scope of constructing models for predicting the output of PV panels, this dataset’s utility extends to other applications. It aids in planning distributed battery energy storage systems aghaei2019flexibility , devising novel energy collection systems kulkarni2018performance , and researching the degradation patterns of photovoltaic panels malvoni2017study . The dataset’s multifaceted applicability emphasizes its significance.
To ensure the validity and precision of the data, careful preprocessing was performed. We discovered approximately 20 anomalies in the original dataset. To maintain a continuous timeline, missing data points were replaced with the arithmetic mean of the preceding and succeeding day’s values. Additionally, data corresponding to the date “12/31/13” was excluded since it contained all-zero values, suggesting an error in the data collection. As a result, we obtained an uninterrupted dataset ranging from 4:00 AM on “3/5/12” to 12:00 AM on “12/30/13.”
Additional in-depth analysis of the dataset was also conducted for a more nuanced understanding. Fig. 2(a) delineates the hourly distribution of PV power across the entire period. As expected, peak PV power values occur during midday, while night time values plummet to zero. Fig. 2(b) portrays monthly PV power fluctuations, which are more volatile compared to daily patterns, likely attributable to the limited number of full-year periods in the dataset. Fig. 2(c) presents a correlation matrix for the dataset features, identifying solar irradiances $I_{3}$ and $I_{15}$ as the features most correlated with PV power. Finally, the joint distribution of dataset features depicted in Fig. 2(c) further confirms that solar intensity is the feature most highly correlated with PV power.
@@ II.2 HQNN
This section introduces our first proposed model, referred to as the HQNN. As illustrated in Fig. 1, the model accepts weather data spanning $24$ consecutive hours as its input. The output is a prediction of the PV power for the upcoming $25$th hour. The HQNN presented in Fig. 3(a) is a combination of classical fully-connected layers, in our case with $120$, $17$ and $8$ neurons, and a VVRQ layer, which is a VQC, consisting of $q$ qubits and $d$ repetitions of variational layers, each distinguished by unique weights. The choice of $120$ neurons is methodical: the model ingests $5$ distinct features for each of the $24$ hours, resulting in a total of $120=5\times 24$. The determination of the remaining parameters stemmed from an extensive hyperparameter optimization process, detailed in the subsequent sections.
@@
Lastly, all the qubits are measured in Pauli-$Z$ basis, yielding the classical vector $v\in\mathbb{R}^{q}$. This output serves as input for a subsequent classical fully-connected layer. This layer processes information from $q$ neurons into $1$ neuron that predicts the power value.
The proposed HQNN model will be compared with its classical analog – a Multilayer Perceptron (MLP) that consists of $4$ fully connected layers with $120$, $32$, $3$, $3$, and $1$ neurons. The number of neurons in each layer was selected by a hyperparameter optimization procedure, detailed in the subsequent sections.
@@ II.3 HQLSTM
This section presents a description of our second hybrid model – HQLSTM, which is a hybrid analog of the classical LSTM model hochreiter1997long , with which predictions will be compared in the following sections. LSTM architectures have garnered significant attention in the realm of time series forecasting, including in predicting PV power gao2023short ; sharadga2020time ; tovar2020pv . HQLSTM models have performed well on tasks such as natural language processing wang2022adverse , detecting software vulnerabilities akter2022automated , and predicting solar radiation yu2023prediction .
In this proposed model, we added a quantum layer to each of the LSTM gates chen2020quantum . Our implementation is depicted in Fig 3(b). The input to the model is:
@@
    The information from the previous step, denoted by a purple circle, $h(t-1)$. It consists of a tensor of size $h_{dim}$. For the initial step, this is simply a zero vector.
These inputs are processed through classical fully-connected layers to yield vectors with a uniform dimension of $4n_{q}$. These vectors are then concatenated through a bitwise addition operation.
@@
The outputs from QDI layers are transformed via classical fully-connected layers to standardize their dimensions to $h_{dim}$. Following this, similar to the classical LSTM, activation functions together with transformations appropriate to each of the four gates, are applied to the outputs originating from the quantum layers. This processing produces the new cell state and hidden state vectors, $C(t)$ and $h(t)$, respectively.
The process operates in a cyclical manner. For each iteration, the vector from the current time step, $x(t)$, and the hidden vector from the previous step, $h(t-1)$, serve as inputs to the HQLSTM. This iterative process is executed as many times as the input width; in our case, the input width equals $24$. Subsequently, all the hidden vectors are concatenated to produce a single composite vector. This vector is then processed through a fully-connected layer consisting of a single neuron, which outputs a value that predicts the PV power.
@@
In order to increase the expressivity, variational layers are positioned both before and after the encoding layer. The first variational layer is illustrated with green rectangles, the encoding layer with blue rectangles, and the final variational layer with purple rectangles.
@@ II.4 HQSeq2Seq
Here we present a hybrid version of the Sequence-to-Sequence (Seq2Seq) model, first introduced in sutskever2014sequence . Seq2seq models are widely used in natural language processing tasks kalchbrenner2013recurrent , where the length of input and output sequence is not predetermined and can be variable. We can also apply the principle of Seq2Seq models to the power prediction task mu2023improved . This means we can feed the neural network with time series of arbitrary length and prompt it to give us the forecast for any number of hours ahead. In this problem setting, the longer the input time series is, the better the model prediction is. Similarly, the shorter the required output length, the easier it is for the model to generate the forecast.
@@
In the QDI layer, instead of attempting to use a qubit for each feature asel1 , we employed the data re-uploading technique schuld2021effect ; data_reuploading . Specifically, we work with four qubits and structure them into a lattice of depth $4$ (depicted as a blue big rectangle in Fig. 3(d)). Each of our $16$ input features leading to the quantum layer is intricately encoded within this lattice. The first four features are mapped onto the initial depth, followed by the subsequent features in blocks of four. Encoding these classical features into the quantum domain, we perform “angle embedding” using $R_{z}$ gates. This operation effectively translates the input vector into a quantum state that symbolizes the preceding classical layer’s data.
@@
Each large blue bounding box encompasses a variational layer (small purple squares). Prior to all the encoding layers, we introduce a single variational layer (small green squares) for enhanced model representation. Consequently, the total weight count in the quantum segment of our network is $20$. In the measurement phase, all the qubits execute a CNOT operation where the first qubit is the target. This ensures that when the first qubit is measured in the $Y$-basis, the measurement outcome depends on the all qubits involved in the circuit. In this way, the output of the quantum layer serves as the power value prediction for a specific hour.
The input size of the encoder and decoder can differ, which is a substantial benefit. For instance, we can use all of the $5$ features to create a context vector, but request to generate the forecast for only $1$ feature. Exploiting this advantage, we will feed the Seq2Seq model with a window of all known features and only demand the power forecast. For the sake of simplicity, we will train both models with a fixed input and output length of $96$ hours and then vary the length in the testing stage.
Figure 4: Results of training and testing the HQNN, MLP, HQLSTM and LSTM models. (a-b) Training and testing history of the models shown by the dotted and solid lines, respectively. The filled space shows the standard deviation of that models averaged over the different testing subsets. (c) Histogram showing MAE and RMSE loss on the test dataset averaged over the results of 5 models trained on different subsets of training and test data. (d) The histogram shows MSE loss on the test dataset for MLP, HQNN, LSTM and HQLSTM models for the reduced training dataset size. (e) The histogram shows the RMSE loss on the test dataset for the Seq2Seq and HQSeq2Seq models for the reduced training dataset size. (f-g) The train and test learning curves for the Seq2Seq and HQSeq2Seq models. (h) An example of the classical and hybrid Seq2Seq models inference on the testing data. The models get $124$ hours of data as an input (before the dashed line) and give a forecast of the “Power” up to $137$ hours ahead (after the dashed line). The solid black line represents the ground truth value of the “Power” feature. The MSE/MAE errors in this particular example are $0.0052/0.0349$ for Seq2Seq and $0.0040/0.0292$ for HQSeq2Seq.
II.5 Training and results
In the study, six distinct models were employed for PV power prediction based on weather features: HQNN, MLP, HQLSTM, LSTM, HQSeq2Seq, and Seq2Seq. To train the models, the mean square error (MSE) was chosen as the loss function:

### v3

@@ Abstract
Accurate forecasting of photovoltaic power is essential for reliable grid integration, yet remains difficult due to highly variable irradiance, complex meteorological drivers, site geography, and device-specific behavior. Although contemporary machine learning has achieved successes, it is not clear that these approaches are optimal: new model classes may further enhance performance and data efficiency. We investigate hybrid quantum neural networks for time-series forecasting of photovoltaic power and introduce two architectures. The first, a Hybrid Quantum Long Short-Term Memory model, reduces mean absolute error and mean squared error by more than $40\%$ relative to the strongest baselines evaluated. The second, a Hybrid Quantum Sequence-to-Sequence model, once trained, it predicts power for arbitrary forecast horizons without requiring prior meteorological inputs and achieves a $16\%$ lower mean absolute error than the best baseline on this task. Both hybrid models maintain superior accuracy when training data are limited, indicating improved data efficiency. These results show that hybrid quantum models address key challenges in photovoltaic power forecasting and offer a practical route to more reliable, data-efficient energy predictions.
quantum machine learning, hybrid quantum neural network, quantum depth-infused layer, quantum LSTM, solar energy, photovoltaic power, time series
@@ I Introduction
Electricity generation prediction, especially for photovoltaic (PV) systems, is a crucial tool for renewable energy adoption sabri2022novel ; li2020review . The global economy must radically reduce emissions to stay within the $1.5\degreeC$ pathway (Paris Agreement) and the transition to renewable energy sources is necessary to achieve these objectives ellabban2014renewable . According to the International Energy Agency, solar PV’s installed power capacity is poised to surpass that of coal by 2027, becoming the world’s largest power source.
@@
Accurate PV power forecasts are vital for multiple facets of the energy industry such as long-term investment planning, regulatory compliance for avoiding penalties, and renewable energy management across storage, transmission, and distribution activities. Several studies show that an increase in forecasting accuracy reduces electricity generation from conventional sources. Increased accuracy also reduces operating costs of systems by reducing the uncertainty of PV power generation anido2016value . Such forecasts support improving the stability and sustainability of the power grid through optimizing power flow and counteracting solar power’s intermittent nature lee2021pv . Such predictions are foundational in increasing the economic viability and improving the adoption of solar energy as they inform pricing and economic dispatch strategies, bolster competitiveness and over time reduce reliance on reserve power. Additionally, they assist with effective energy-storage management and the integration of PV systems into the power grid bozorgavari2019two , which is essential for the enduring success of renewable energy solutions diagne2013review .
Traditional methods for predicting PV power have primarily relied on statistical models, machine learning algorithms, or a blend of both ahmed2019review . In Ref. sabri2022novel , a hybrid GRU–CNN model is proposed for forecasting the output of the DKASC 1B PV system (Alice Springs), showing improvements across all metrics compared with standalone GRU/LSTM/CNN models. In Ref. gomez2020photovoltaic , the authors test an MLP in three scenarios for a plant in Monteroni di Lecce (2012–2013) and obtain nRMSE values of $2.94–9.98\%$. Another study kaloop2021novel involves the training ANFIS-based hybrid algorithms for the University of Salento PV plant, achieving a minimum RMSE of $0.081$. Applying kNN to the University of Salento data produced an absolute error of $5.5\%$. In Ref. shi2012forecasting , the time series of a grid-connected PV plant in southern China are forecast using SVM models, yielding an average MRE of $8.64\%$. A recent study implements a hybrid “FFNN + SVM” scheme optimized with particle swarm optimization (PSO) for long-term forecasting using real climate and load data; it achieves very high accuracy (correlation coefficient $R^{2}\approx 0.9984$) but notes that random variability in the inputs can slow model convergence boum2022photovoltaic . Another work proposes a deep-learning ensemble—an MLP combined with an LSTM—which outperforms classical approaches on demand and generation datasets mbey2024novel ; in a subsequent study, a multi-objective PSO is integrated into an FFNN–LSTM hybrid, further reducing errors (MAPE $0.1\%$) for joint PV and load forecasting at the cost of increased model complexity mbey2024solar . In addition, an ensemble of SVM and GRU models optimized by an ant colony algorithm aggregates their outputs and captures weather-induced variability across multiple PV sites, achieving a high level of agreement (correlation coefficient $R^{2}\approx 0.9986$souhe2024optimized , though the authors emphasize the need for more advanced fusion strategies and real-time adaptation for practical deployment. Nevertheless, the variable and nonlinear nature of solar generation, influenced by a wide range of meteorological factors, remains a serious challenge for traditional models blaga2019current .
@@
In light of these challenges, quantum machine learning (QML) emerges as a promising avenue. This rapidly evolving field, which melds the principles of quantum mechanics with classical machine learning Schuld_2014 ; biamonte2017quantum ; Dunjko.Briegel.2018 ; qml_review_2023 , can offer enhanced capabilities for improving the forecasting accuracy of time series tasks emmanoulopoulos2022quantum , including PV power generation sushmit2023forecasting . QML’s potential arises from quantum features like superposition and entanglement, promising exponential speedups in certain tasks schuld2018implementing . Moreover, QML algorithms produce inherently probabilistic results, aptly suited for prediction tasks, and they may potentially function within an exponentially larger search space, amplifying their efficacy Lloyd2013 ; 365700 ; Lloyd1996UniversalQS ; schuld2020circuit . Nonetheless, implementing quantum algorithms bears its own set of challenges, such as the need for error correction and sensitivity to external interference Aaronson2015 . Yet, in spite of these challenges, hybrid quantum-classical models, especially hybrid quantum neural networks (HQNNs), have showcased their potential in diverse industrial realms, including healthcare jain2022hybrid ; pharma ; steatosis_sagingalieva ; sagingalieva2025hybrid ; anoshin2024hybrid , energy kurkin2025forecasting ; sushmit2023forecasting ; lee2025predictive , aerospace rainjonneau2023quantum , logistics haboury2023supervised and, automotive sagingalieva2023hyperparameter .
@@
Figure 1: The input to a hybrid quantum model is presented as a chronological data table, documenting hourly meteorological parameters including ambient temperature ($T_{a}$), module temperature ($T_{m}$), and solar irradiance ($I_{3}$, $I_{15}$), alongside the mean PV power output ($P$). The model is designed to leverage these data to generate predictions of PV power output for a short-term forecast aimed at near-future output, typically the next hour, and a long-term forecast that extends to a broader temporal horizon.
@@
In this article, we present three types of hybrid quantum models as potential solutions for PV power forecasting. We assess the performance of our proposed models using a publicly accessible dataset malvoni2016data , encompassing a comprehensive array of meteorological variables as well as hourly mean PV power measurements spanning a $21$-month period. This dataset, along with the data preprocessing and analytical methods employed, is described in detail in Section II.1.
@@
Despite using fewer trainable parameters, our hybrid quantum models outperform their classical counterparts in terms of more accurate predictions, including when trained on a reduced dataset. We summarize our conclusions and outline future research directions in Section III.
II Materials and methods
@@
The entire training workflow can be summarized as a five-stage pipeline, each stage being detailed in subsequent subsections:
  1. 1.
    Data preprocessing.
    1. (a)
      A systematic cleaning pass is performed in which missing and inconsistent records are addressed. Gaps are imputed via piecewise–linear interpolation. An exploratory dependency analysis is conducted to identify dominant correlations among meteorological variables and PV output. Finally, all numeric features are rescaled to the unit interval $[0,1]$ using statistics computed only on the training partition to avoid target leakage; the same transformation is applied thereafter to validation and test splits.
    2. (b)
      After being reduced to a common format, the series is transformed into a supervised learning format based on sliding windows tailored to the forecasting task. Each sample is defined by three parameters: the input window length (history used as context), the output window length (forecast horizon), and the stride by which the shift between consecutive windows is set. Multi–horizon prediction is supported, and temporal ordering preserved by chronological splits.
  2. 2.
    Model development. Selection of the overall model architecture, including the choice between hybrid and classical configurations, the number of hidden layers, the types of layers (recurrent or feed-forward, quantum or classical), and the activation functions.
  3. 3.
    Analysis of quantum layers. For hybrid models, the embedded quantum circuit is analyzed to identify configurations that are able to maintain trainability and expressivity balance. Structural reducibility and over-parameterization are examined, and ranges for circuit hyperparameters are chosen to avoid barren–plateau behavior.
  4. 4.
    Hyperparameter optimization and training. Hyperparameter optimization for both classical and hybrid models is performed using specialized frameworks, with a search space including learning rate, layer size, dropout rate, number of variational layers, and number of qubits. The optimal configuration is then trained using $k$-fold cross-validation.
  5. 5.
    Model performance. Performance is assessed and then aggregated across folds using a comprehensive set of metrics, including MAE, MSE, RMSE, MAPE, $R^{2}$, and VAF. A final averaged estimate is provided together with a discussion of trade–offs between accuracy and efficiency.
@@ II.1 Dataset
To investigate the advantages of hybrid quantum models, we selected a publicly accessible dataset malvoni2016data from a conventional generation plant situated in the Mediterranean region. This dataset is not only comprehensive but also enables benchmarking with results from various algorithms available in the literature. A comparative analysis of our model’s predictions and those from the study by Ref. kaloop2021novel is provided in Section III. The dataset, presented as a numerical table showcased in Fig. 1, includes variables such as hourly mean ambient temperature, $T_{a}$; hourly mean module temperature, $T_{m}$; hourly mean solar irradiance recorded on two tilted planes with tilt angles of 3 and 15 degrees, $I_{3}$ and $I_{15}$; and hourly mean PV power, $P$, spanning $21$ months, accounting for more than $500$ days.
Beyond the scope of constructing models for predicting the output of PV panels, this dataset’s utility extends to other applications. It aids in planning distributed battery energy storage systems aghaei2019flexibility , devising novel energy collection systems kulkarni2018performance , and researching the degradation patterns of PV panels malvoni2017study . The dataset’s multifaceted applicability emphasizes its significance.
To ensure the validity and precision of the data, careful preprocessing was performed. We discovered approximately 20 anomalies in the original dataset. To maintain a continuous timeline, missing data points were replaced with the arithmetic mean of the preceding and succeeding days values. Additionally, data corresponding to the date “12/31/13” were excluded since it contained all-zero values, suggesting an error in the data collection. As a result, we obtained an uninterrupted dataset ranging from 4:00 AM on “3/5/12” to 12:00 AM on “12/30/13.”
Additional in-depth analysis of the dataset was also conducted for a more nuanced understanding. Fig. 2(a) delineates the hourly distribution of PV power across the entire period. As expected, peak PV power values occur during midday, while nighttime values plummet to zero. Fig. 2(b) portrays monthly PV power fluctuations, which are more volatile compared to daily patterns, likely attributable to the limited number of full-year periods in the dataset. Fig. 2(c) presents a correlation matrix for the dataset features, identifying solar irradiances $I_{3}$ and $I_{15}$ as the features most correlated with PV power. Finally, the joint distribution of dataset features depicted in Fig. 2(d) further confirms that solar intensity is the feature most highly correlated with PV power.
@@ II.2 HQNN
This section introduces our first proposed model, referred to as the HQNN. As illustrated in Fig. 1, the model accepts weather data spanning $24$ consecutive hours as its input. The output is a prediction of the PV power for the upcoming $25$th hour. The HQNN presented in Fig. 3(a) is a combination of classical fully-connected layers, in our case with $120$, $17$, and $8$ neurons, and a VVRQ layer, which is a VQC, consisting of $q$ qubits and $d$ repetitions of variational layers, each distinguished by unique weights. The choice of $120$ neurons is methodical: the model ingests $5$ distinct features for each of the $24$ hours, resulting in a total of $120=5\times 24$. The determination of the remaining parameters stemmed from an extensive hyperparameter optimization process, detailed in the subsequent sections.
@@
Lastly, all the qubits are measured in the Pauli-$Z$ basis, yielding the classical vector $v\in\mathbb{R}^{q}$. This output serves as input for a subsequent classical fully-connected layer. This layer processes information from $q$ neurons into $1$ neuron that predicts the power value.
The proposed HQNN model will be compared with its classical analogue – a MLP that consists of $5$ fully connected layers with $120$, $32$, 33, 33, and $1$ neurons. The number of neurons in each layer was selected by a hyperparameter optimization procedure, detailed in the subsequent sections.
@@ II.3 HQLSTM
This section presents a description of our second hybrid model – HQLSTM, which is a hybrid analogue of the classical LSTM model hochreiter1997long , with which predictions will be compared in the following sections. LSTM architectures have garnered significant attention in the realm of time series forecasting, including in predicting PV power gao2023short ; sharadga2020time ; tovar2020pv . HQLSTM models have performed well on tasks such as natural language processing wang2022adverse , detecting software vulnerabilities akter2022automated , and predicting solar radiation yu2023prediction .
In this proposed model, we added a quantum layer to each of the LSTM gates chen2020quantum . Our implementation is depicted in Fig. 3(b). The input to the model is:
@@
    The information from the previous step, denoted by a purple circle, h(t1)h(t-1). It consists of a tensor of size $h_{dim}$. For the initial step, this is simply a zero vector.
These inputs are processed through classical fully-connected layers to yield vectors with a uniform dimension of $4n_{q}$. These vectors are then combined via element-wise addition.
@@
The outputs from QDI layers are transformed via classical fully-connected layers to standardize their dimensions to $h_{dim}$. Following this, similar to the classical LSTM, activation functions together with transformations appropriate to each of the four gates are applied to the outputs originating from the quantum layers. This processing produces the new cell state and hidden state vectors, $C(t)$ and $h(t)$, respectively.
The process operates in a cyclical manner. For each iteration, the vector from the current time step, $x(t)$, and the hidden vector from the previous step, h(t1)h(t-1), serve as inputs to the HQLSTM. This iterative process is executed as many times as the input width; in our case, the input width equals $24$. Subsequently, all the hidden vectors are concatenated to produce a single composite vector. This vector is then processed through a fully-connected layer consisting of a single neuron, which outputs a value that predicts the PV power.
@@
To increase the expressivity, variational layers are positioned both before and after the encoding layer. The first variational layer is illustrated with green rectangles, the encoding layer with blue rectangles, and the final variational layer with purple rectangles.
@@ II.4 HQSeq2Seq
Here we present a hybrid version of the Sequence-to-Sequence (Seq2Seq) model, first introduced in Ref. sutskever2014sequence . Seq2Seq models are widely used in natural language processing tasks kalchbrenner2013recurrent , where the length of input and output sequences is not predetermined and can be variable. We can also apply the principle of Seq2Seq models to the power prediction task mu2023improved . This means we can feed the neural network with time series of arbitrary length and produce a forecast for any number of hours ahead. In this problem setting, the longer the input time series is, the better the model prediction is. Similarly, the shorter the required output length, the easier it is for the model to generate the forecast.
@@
In the QDI layer, instead of attempting to use a qubit for each feature asel1 , we employed the data re-uploading technique schuld2021effect ; data_reuploading . Specifically, we work with four qubits and structure them into a lattice of depth $4$ (depicted as a large blue rectangle in Fig. 3(d)). Each of our $16$ input features leading to the quantum layer are intricately encoded within this lattice. The first four features are mapped onto the initial depth, followed by the subsequent features in blocks of four. Encoding these classical features into the quantum domain, we perform “angle embedding” using $R_{z}$ gates. This operation effectively translates the input vector into a quantum state that symbolizes the preceding classical layer’s data.
@@
Each large blue bounding box encompasses a variational layer (small purple squares). Prior to all the encoding layers, we introduce a single variational layer (small green squares) for enhanced model representation. Consequently, the total weight count in the quantum segment of our network is $20$. In the measurement phase, all the qubits execute a CNOT operation where the first qubit is the target. This ensures that when the first qubit is measured in the $Y$-basis, the measurement outcome depends on all the qubits involved in the circuit. In this way, the output of the quantum layer serves as the predicted power value for a specific hour.
The input size of the encoder and decoder can differ, which is a substantial benefit. For instance, we can use all of the $5$ features to create a context vector, but request to generate the forecast for only $1$ feature. Exploiting this advantage, we will feed the Seq2Seq model with a window of all known features and only demand the power forecast. For the sake of simplicity, we trained both models with a fixed input and output length of $96$ hours and then vary the length in the testing stage.
Figure 4: Results of training and testing for the HQNN, MLP, HQLSTM, and LSTM models. (a-b) Training and testing history shown by the dotted and solid lines, respectively. The filled region shows the standard deviation of the models averaged over the different testing subsets. (c) Bar chart showing MAE and RMSE loss on the test dataset averaged over the results of 5 models trained on different subsets of training and test data. (d) The bar chart shows MSE loss on the test dataset for MLP, HQNN, LSTM and HQLSTM models for the reduced training dataset size. (e) The bar chart shows the RMSE loss on the test dataset for the Seq2Seq and HQSeq2Seq models for the reduced training dataset size. (f-g) The training and testing learning curves for the Seq2Seq and HQSeq2Seq models. (h) Example for the classical and hybrid Seq2Seq models inference on the testing data. The models receive $124$ hours of data as an input (before the dashed line) and forecast the “Power” up to $137$ hours ahead (after the dashed line). The solid black line represents the ground-truth value of the “Power” feature. The MSE/MAE errors in this particular example are $0.0052/0.0349$ for Seq2Seq and $0.0040/0.0292$ for HQSeq2Seq.
III Results and discussion
In the study, six distinct models were employed for PV power prediction based on weather features: HQNN, MLP, HQLSTM, LSTM, HQSeq2Seq, and Seq2Seq. To train the models, the mean squared error (MSE) was chosen as the loss function: